# Constrained Denoising, Empirical Bayes, and Optimal Transport

### v1

@@ 3 Oracle Constrained Bayes Denoising
In this section, we consider the oracle versions (i.e., assuming that the latent variable distribution $G$ is known) of our constrained denoising problems ($\mathcal{VCB}$), ($\mathcal{DCB}$) and ($\mathcal{GCB}$). We address the case of variance constraints in Section 3.1, the case of distributional constraints in Section 3.2, and the case of general constraints in Section 3.3. The proofs of all the results stated in this section are given in Appendix D.
@@ 3.1 Variance-Constrained Denoising
As mentioned in Section 1, it is a common practice across scientific domains to plot the empirical distribution of Bayes denoisers $\delta_{\mathcal{B}}(Z_{1}),\dotsc,\delta_{\mathcal{B}}(Z_{n})$ or of EB estimates thereof. Example areas include astronomy (Eddington, 1940; Loredo, 2007) and multisite trials in education research (Raudenbush and Bloom, 2015; Bloom et al., 2017; Lee et al., 2024). Graphical distributional summaries (e.g., histograms, density plots) of Bayes denoisers inevitably exhibit underdispersion relative to the true distribution of the unobserved latent variables $\Theta_{1},\dotsc,\Theta_{n}$ (as directly follows from (1.2)). At an intuitive level, these graphical summaries would more faithfully represent the true heterogeneity in the population if the denoisers could be “unshrunk” to match the variance of the latent variables. Problem ($\mathcal{VCB}$) captures precisely this intuition through its variance-matching constraint, as Theorem 3.3 below demonstrates. Thus, problem ($\mathcal{VCB}$) provides a principled and useful strategy for presenting more faithful empirical summaries of heterogeneity by at least capturing the correct second moment structure. We note that this motivation was central to the work of Louis (1984) and Ghosh (1992).
@@
Consider a regression setting with covariates $X\in\mathbb{R}^{k}$ and response $Y\in\mathbb{R}$; interest focuses on the relationship between $X$ and $Y$. We observe the covariates with noise, that is, we observe $W=X+E$ where $E$ has a known distribution. In regression calibration (Carroll and Stefanski, 1990) we perform the regression analysis on the pairs $(\mathbb{E}[X\mid W],Y)$. By contrast, in moment reconstruction (Freedman et al., 2004), we seek a denoiser of $X$, which may possibly depend on both $W$ and $Y$, that satisfies
@@
Our analysis of problem ($\mathcal{VCB}$) requires a few assumptions, which we now introduce and discuss. Because we will impose covariance constraints on the denoising problem, we need some conditions to ensure that the constrained denoising problem admits a solution. For example, the following is necessary:
@@
Note that we require the above second moment condition for the distribution $G$ but not for $F$ (the marginal distribution of $Z$). We also make the following assumption on the covariance of the Bayes denoiser $\delta_{\mathcal{B}}$:
@@
We are now ready to state the first of our main results, for which the univariate case is effectively due to Louis (1984); Ghosh (1992). Recall that $R_{\mathcal{B}}$ represents the minimum value of problem ($\mathcal{B}$) and that $R_{\mathcal{VCB}}$ represents the minimum value of problem ($\mathcal{VCB}$). Also, recall the definition of $\mathbf{t}$ given in (2.3).
@@
Under assumption (2M) and (PD), problem ($\mathcal{VCB}$) has a solution $\delta_{\mathcal{VCB}}:\mathbb{R}^{d}\to\mathbb{R}^{m}$ given by
@@
and its risk is given by
$R_{\mathcal{VCB}}=R_{\mathcal{B}}+W_{2}^{2}\bigg{(}\mathcal{N}(0,\textnormal{Cov}(\Theta)),\;\;\mathcal{N}(0,\textnormal{Cov}(\delta_{\mathcal{B}}(Z)))\bigg{)}.$ (3.2)
Under the further assumptions (ZC) and (BC), $\delta_{\mathcal{VCB}}$ is the unique solution to problem ($\mathcal{VCB}$), up to $F$-almost everywhere equivalence.
@@
The multivariate problem of optimal denoising subject to constraints on the covariance matrix has been previously studied by Ghosh and Maiti (1999), and their proposed denoiser is
$\delta(z)=(\textnormal{Cov}(\Theta))^{\nicefrac{{1}}{{2}}}(\textnormal{Cov}(\delta_{\mathcal{B}}(Z)))^{-\nicefrac{{1}}{{2}}}(\delta_{\mathcal{B}}(z)-\mathbb{E}[\Theta])+\mathbb{E}[\Theta].$
Note that this differs from the optimal denoiser given in Theorem 3.3, but both are of the form
@@
for a suitable matrix $T\in\mathbb{R}^{m\times m}$; that is, both are an affine transformation of the Bayes denoiser. We reconcile this difference by observing that an additional condition is needed in the discussion following Ghosh and Maiti (1999, equation (40)): Ghosh and Maiti (1999) identify that an optimal denoiser must be of the form (3.3), but omit the condition that $T$ be symmetric. Since all subsequent applications (e.g., to the dataset in Devine et al. (1994)) involve matrices $\textnormal{Cov}(\Theta)$ and $\textnormal{Cov}(\delta_{\mathcal{B}}(Z))$ that commute, it holds that
$\mathbf{t}_{\textnormal{Cov}(\delta_{\mathcal{B}}(Z))}^{\textnormal{Cov}(\Theta)}=(\textnormal{Cov}(\Theta))^{\nicefrac{{1}}{{2}}}(\textnormal{Cov}(\delta_{\mathcal{B}}(Z)))^{-\nicefrac{{1}}{{2}}},$
and so, in this case, the denoiser in Ghosh and Maiti (1999) is identical to the denoiser of Theorem 3.3. Stating the result in the general case when $\textnormal{Cov}(\Theta)$ and $\textnormal{Cov}(\delta_{\mathcal{B}}(Z))$ do not commute requires the machinery of the Bures-Wasserstein geometry as summarized in Section 2.3.
At this point, a natural goal is to quantify the statistical cost of imposing the moment matching constraint. More precisely, one may ask how large is the excess risk $R_{\mathcal{VCB}}-R_{\mathcal{B}}$. Our next result (proved in Appendix D) shows that the price to pay is modest.
@@ 3.2 Distribution-Constrained Denoising
In Example 3.2, we introduced moment reconstruction. Freedman et al. (2004) propose moment reconstruction with the following motivation. Suppose we can come up with a denoiser $\delta(W,Y)$ of $X$ such that,
[…]
then we could learn about the relationship between $(X,Y)$ by studying $(\delta(W,Y),Y)$. Note that the constraint above corresponds to the distribution-constrained Bayes problem in ($\mathcal{DCB}$). Freedman et al. (2004) instead propose the moment reconstruction approach described in Example 3.2, arguing that enforcing the full distributional constraint “is a very difficult problem, but if we content ourselves with the lesser aim of matching just the first two moments of the joint distribution, then a simple solution can be obtained.”
@@
which, as we emphasize in the notation, depends on the unknown distribution $G$ (and also the known likelihood $P=\{P_{\theta}\}_{\theta}$) via the conditional expectation $\delta_{\mathcal{B}}(z)=\mathbb{E}[\Theta\,|\,Z=z]$. We then have the following, which is a re-statement of García-Trillos and Sen (2024, Theorem 2.4) and bears some similarity to Freirich et al. (2021, Theorem 1). Recall from Section 2 that $\Gamma(F;G)$ represents the space of all couplings of $F$ and $G$.
@@
$\begin{cases}\textnormal{minimize}&\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}(z,\eta)\,\textnormal{d}\pi(z,\eta)\\ \textnormal{over}&\pi\in\Gamma(F;G),\end{cases}$ (3.5)
has a unique solution $\pi_{\mathcal{DCB}}$. This solution is concentrated on the graph of a function which we denote by $\delta_{\mathcal{DCB}}:\mathbb{R}^{d}\to\mathbb{R}^{m}$. This function $\delta_{\mathcal{DCB}}$ can be written as $\delta_{\mathcal{DCB}}=\nabla\phi\circ\delta_{\mathcal{B}}$ for some convex function $\phi:\mathbb{R}^{m}\to\mathbb{R}\cup\{+\infty\}$, and we have
$R_{\mathcal{DCB}}=R_{\mathcal{B}}+W_{2}^{2}(G,(\delta_{\mathcal{B}})_{\#}F).$
Consequently, (3.5) is a tight convex relaxation of ($\mathcal{DCB}$), and $\delta_{\mathcal{DCB}}$ is the unique solution to ($\mathcal{DCB}$) up to $F$-almost everywhere equivalence.
The above result says that, under appropriate conditions, problem ($\mathcal{DCB}$) indeed has a unique solution which can be obtained by solving the linear program (3.5). Further, the risk of the distribution-constrained denoiser $\delta_{\mathcal{DCB}}(Z)$ is the sum of the Bayes risk $R_{\mathcal{B}}$ and the squared Wasserstein distance between the distributions of $\Theta$ and $\delta_{\mathcal{B}}(Z)$. Contrast this with (3.2), where we showed that the risk of the variance-constrained denoiser $\delta_{\mathcal{VCB}}(Z)$ is sum of the Bayes risk and the squared Bures-Wasserstein distance between the covariance matrices of $\Theta$ and $\delta_{\mathcal{B}}(Z)$.
@@ 3.3 General-Constrained Denoising
Lastly, we consider problem ($\mathcal{GCB}$) of general-constrained Bayes denoising. That is, we fix a linearly-independent collection $\psi_{1},\ldots,\psi_{k}:\mathbb{R}^{m}\to\mathbb{R}$ of continuous functions (which are also assumed to be $G$-integrable). For convenience, we write $\Psi:=\{\psi_{1},\ldots,\psi_{k}\}$. Roughly speaking, any constraint on the distribution of $\delta(Z)$ can be cast in the form above (although we may need to take $k$ to be infinite); indeed, ($\mathcal{GCB}$) reduces to ($\mathcal{B}$) if $\Psi$ is empty, ($\mathcal{VCB}$) if $\Psi$ consists of all quadratic forms, and ($\mathcal{DCB}$) if $\Psi$ consists of all bounded continuous functions. In addition to subsuming these earlier examples, we are motivated by the following.
Example 3.9 (Higher moment constraints).
$\begin{cases}\textnormal{minimize}&\mathbb{E}\left[|\delta(Z)-\Theta|^{2}\right]\\ \textnormal{over}&\delta:\mathbb{R}\to\mathbb{R}\\ \textnormal{with}&\mathbb{E}\left[(\delta(Z))^{\ell}\right]=\mathbb{E}\left[\Theta^{\ell}\right]\quad\textnormal{ for all }\ell=1,\ldots,k.\end{cases}$ (3.6)
This, of course, coincides with the case of variance constraints when $k=2$, and higher moment constraints force the distribution of $\delta(Z)$ to be even closer to the distribution $G$ of $\Theta$. Along these lines it is shown in Armstrong et al. (2022), in the context of constructing nonparametric EB confidence intervals, that imposing constraints on the fourth moment leads to EB procedures which nearly adapt to the parametric (Gaussian) setting.
Example 3.10 (Errors-in-variables regression continued: moment adjusted imputation).
Bay (1997) and Thomas et al. (2011) propose an intermediate approach called moment adjusted imputation that bridges the methods described in Examples 3.2 and 3.7. Their approach uses a denoiser that matches a controlled number of moments and cross-moments between $X$ and $Y$, extending beyond the two moments in moment reconstruction but without enforcing full distributional equivalence. For instance, when $X\in\mathbb{R}$, they aim to estimate $X$ while preserving multiple moments such as $\mathbb{E}[X]$, $\mathbb{E}[X^{2}]$, $\mathbb{E}[X^{3}]$, $\mathbb{E}[X^{4}]$, as well as cross-moments like $\mathbb{E}[XY]$ and $\mathbb{E}[X^{2}Y]$. Such a denoiser is given by problem ($\mathcal{GCB}$) with the choices $Z=(W,Y)$, $\Theta=(X,Y)$, $\psi_{\ell}(x,y)=x^{\ell}$ for $\ell=1,\dotsc,4$, and $\psi_{5}(x,y)=xy$, $\psi_{6}(x,y)=x^{2}y$.
@@
Suppose that $G$ is known to have support contained in a closed set $S\subseteq\mathbb{R}^{m}$. In that case, the posterior mean satisfies $\delta_{\mathcal{B}}(Z)\in S$ almost surely. However, the variance constrained denoiser $\delta_{\mathcal{VCB}}(\cdot)$ may take on values in $S$ with positive probability, which can be undesirable. Instead, one may aim to minimize the denoising risk subject to simultaneous constraints on the variance and the support of $\delta(Z)$, thereby solving the problem
$\begin{cases}\textnormal{minimize}&\mathbb{E}\left[\|\delta(Z)-\Theta\|^{2}\right]\\ \textnormal{over}&\delta:\mathbb{R}^{d}\to\mathbb{R}^{m}\\ \textnormal{with}&\textnormal{Cov}(\delta(Z))=\textnormal{Cov}(\Theta)\\ \textnormal{and}&\mathbb{P}(\delta(Z)\in S)=1.\end{cases}$ (3.7)
To see that this can be cast as a special case of problem ($\mathcal{GCB}$), simply note that the constraint $\mathbb{P}(\delta(Z)\in S)=1$ can be equivalently written as $\mathbb{E}[\min_{\theta\in S}\|\delta(Z)-\theta\|^{2}]=0$. (We can replace $\psi(\eta):=\min_{\theta\in S}\|\eta-\theta\|$ with any function $\psi:\mathbb{R}^{m}\to[0,\infty)$ satisfying $\psi(\eta)>0$ if and only if $\eta\notin S$.) In the special case of $m=d=1$ an important constraint on the support is nonnegativity, i.e., $S=[0,\infty)$.
Furthermore, we assume that
[…]
We note that (C) is a mild condition, and that it is satisfied whenever $\{P_{\theta}\}_{\theta}$ is a suitable exponential family, regardless of $G$ (García-Trillos and Sen, 2024, Remark 2.2); for example, it holds for Gaussian or Poisson families. (It holds trivially whenever $F$ has discrete support.) Lastly, we assume that
[…]
Theorem 3.12.
Under assumptions (ZC), (BC), (C), and (QG), the problem
$\begin{cases}\textnormal{minimize}&\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}(z,\eta)\,\textnormal{d}\pi(z,\eta)\\ \textnormal{over}&\pi\in\Gamma(F;G,\Psi),\end{cases}$ (3.8)
has a unique solution $\pi_{\mathcal{GCB}}$, and this solution is concentrated on the graph of a function which we denote $\delta_{\mathcal{GCB}}:\mathbb{R}^{d}\to\mathbb{R}^{m}$. If $H$ denotes the marginal distribution of $D$ when $(Z,D)\sim\pi_{\mathcal{GCB}}$, then $\delta_{\mathcal{GCB}}$ can be written as $\delta_{\mathcal{GCB}}=\nabla\phi\circ\delta_{\mathcal{B},H}$ for some convex function $\phi:\mathbb{R}^{m}\to\mathbb{R}\cup\{+\infty\}$, and we have
$R_{\mathcal{GCB}}=R_{\mathcal{B}}+W_{2}^{2}(H,(\delta_{\mathcal{B}})_{\#}F).$
Consequently, (3.8) is a tight convex relaxation of ($\mathcal{GCB}$), and $\delta_{\mathcal{GCB}}$ is the unique solution to ($\mathcal{GCB}$) up to $F$-almost everywhere equivalence.
As in the case of problem ($\mathcal{DCB}$), this result provides conditions under which problem ($\mathcal{GCB}$) has a unique solution which can be found by solving the linear program (3.8). We refer to the probability measure as $H$ as the projected latent variable distribution since, among all probability measures whose integrals with $\Psi$ match those of $G$, it is the nearest to $(\delta_{\mathcal{B}})_{\#}F$ with respect to the distance $W_{2}$.
Example 3.13 (Higher moment constraints continued).
To provide further intuition on the projected latent variable distribution $H$, we characterize it in some particular cases of Example 3.9. If $k=0$ or $k=1$, then $\delta_{\mathcal{B}}$ is feasible hence optimal for (3.6), so $H$ equals $(\delta_{\mathcal{B}})_{\#}F$. If $k=2$, then (3.6) is equivalent to ($\mathcal{VCB}$), so $H$ equals the pushforward of $F$ by an affine function of $\delta_{\mathcal{B}}$. If $G$ is compactly supported and $H_{k}$ denotes the projected latent variable distribution for constraints on $k$ moments, then $H_{k}$ converges weakly to $G$ as $k\to\infty$.

### v2

@@ 3 Oracle Constrained Bayes Denoising
In this section, we consider the oracle versions (i.e., assuming that the latent variable distribution $G$ is known) of our constrained denoising problems ($\mathcal{VCB}$), ($\mathcal{DCB}$) and ($\mathcal{GCB}$). We address the case of variance constraints in Section 3.1, the case of distributional constraints in Section 3.2, and the case of general constraints in Section 3.3. The proofs of all the results stated in this section are given in Appendix F.
@@ 3.1 Variance-Constrained Denoising
Graphical distributional summaries (e.g., histograms, density plots) of Bayes denoisers inevitably exhibit underdispersion relative to the true distribution of the unobserved latent variables $\Theta_{1},\dotsc,\Theta_{n}$ (as directly follows from (1.2)). At an intuitive level, these graphical summaries would more faithfully represent the true heterogeneity in the population if the denoisers could be “unshrunk” to match the variance of the latent variables. Problem ($\mathcal{VCB}$) captures precisely this intuition through its variance-matching constraint, as Theorem 3.3 below demonstrates. Thus, problem ($\mathcal{VCB}$) provides a principled and useful strategy for presenting more faithful empirical summaries of heterogeneity by at least capturing the correct second moment structure. We note that this motivation was central to the work of Louis (1984) and Ghosh (1992).
@@
Consider a regression setting with covariates $X\in\mathbb{R}^{k}$ and response $Y\in\mathbb{R}$; interest focuses on the relationship between $X$ and $Y$. We observe the covariates with noise, that is, we observe $W=X+E$ where $E$ has a known distribution. In regression calibration (Carroll and Stefanski, 1990) we perform the regression analysis on the pairs $(\mathbb{E}[X\mid W],Y)$; see Chen et al. (2025) for a critical discussion of the approach. By contrast, in moment reconstruction (Freedman et al., 2004), we seek a denoiser of $X$, which may possibly depend on both $W$ and $Y$, that satisfies
@@
Our analysis of problem ($\mathcal{VCB}$) requires a few assumptions, which we now introduce and discuss. Because we will impose covariance constraints on the denoising problem, we need some conditions to ensure that the constrained denoising problem admits a solution. The following is necessary:
@@
We require the second moment condition for $G$ but not for $F$ (the marginal distribution of $Z$). We also make the following assumption on the covariance of the Bayes denoiser $\delta_{\mathcal{B}}$:
@@
We are ready to state the first of our main results, for which the univariate case is effectively due to Louis (1984); Ghosh (1992). Here, $R_{\mathcal{B}}$ denotes the minimum value of problem ($\mathcal{B}$) and $R_{\mathcal{VCB}}$ the minimum value of problem ($\mathcal{VCB}$). Also, recall the definition of $\mathbf{t}$ given in Section 2.
@@
Under assumption (2M) and (PD), problem ($\mathcal{VCB}$) has solution
@@
and its risk is given by $R_{\mathcal{VCB}}=R_{\mathcal{B}}+W_{2}^{2}(\mathcal{N}(0,\textnormal{Cov}(\Theta)),\mathcal{N}(0,\textnormal{Cov}(\delta_{\mathcal{B}}(Z)))$. Under the further assumptions (ZC) and (BC), $\delta_{\mathcal{VCB}}$ is the $F$-a.s. unique solution to problem ($\mathcal{VCB}$).
@@
The multivariate problem of optimal denoising subject to constraints on the covariance matrix has been previously studied by Ghosh and Maiti (1999). Their proposed denoiser is $\delta(z)=(\textnormal{Cov}(\Theta))^{\nicefrac{{1}}{{2}}}(\textnormal{Cov}(\delta_{\mathcal{B}}(Z)))^{-\nicefrac{{1}}{{2}}}(\delta_{\mathcal{B}}(z)-\mathbb{E}[\Theta])+\mathbb{E}[\Theta]$. Note that this differs from the optimal denoiser given in Theorem 3.3, but both are of the form
@@
for a suitable matrix $T\in\mathbb{R}^{m\times m}$; that is, both are an affine transformation of the Bayes denoiser. We reconcile this difference by observing that an additional condition is needed in the discussion following Ghosh and Maiti (1999, equation (40)): Ghosh and Maiti (1999) identify that an optimal denoiser must be of the form (3.2), but omit the condition that $T$ be symmetric. Since all subsequent applications (e.g., to the dataset in Devine et al. (1994)) involve matrices $\textnormal{Cov}(\Theta)$ and $\textnormal{Cov}(\delta_{\mathcal{B}}(Z))$ that commute, it holds that $\mathbf{t}_{\textnormal{Cov}(\delta_{\mathcal{B}}(Z))}^{\textnormal{Cov}(\Theta)}=(\textnormal{Cov}(\Theta))^{\nicefrac{{1}}{{2}}}(\textnormal{Cov}(\delta_{\mathcal{B}}(Z)))^{-\nicefrac{{1}}{{2}}},$ and so, in this case, the denoiser in Ghosh and Maiti (1999) is identical to the denoiser of Theorem 3.3. Stating the result in the general case when $\textnormal{Cov}(\Theta)$ and $\textnormal{Cov}(\delta_{\mathcal{B}}(Z))$ do not commute requires the machinery of the Bures-Wasserstein geometry as summarized in Section 2.
At this point, a natural goal is to quantify the statistical cost of imposing the moment matching constraint. More precisely, one may ask how large is the excess risk $R_{\mathcal{VCB}}-R_{\mathcal{B}}$. Our next result (proved in Appendix F) shows that the price to pay is modest.
@@ 3.2 Distribution-Constrained Denoising
In Example 3.2, we introduced moment reconstruction. Freedman et al. (2004) propose moment reconstruction with the following motivation. Suppose we can come up with a denoiser $\delta(W,Y)$ of $X$ such that, $(\delta(W,Y),Y)\overset{\mathcal{D}}{=}(X,Y),$ then we could learn about the relationship between $(X,Y)$ by studying $(\delta(W,Y),Y)$. Note that the constraint above corresponds to the distribution-constrained Bayes problem in ($\mathcal{DCB}$). Freedman et al. (2004) instead propose the moment reconstruction approach described in Example 3.2, arguing that enforcing the full distributional constraint “is a very difficult problem, but if we content ourselves with the lesser aim of matching just the first two moments of the joint distribution, then a simple solution can be obtained.”
@@
which, as we emphasize in the notation, depends on the unknown distribution $G$ (and also the known likelihood $P=\{P_{\theta}\}_{\theta}$) via the conditional expectation $\delta_{\mathcal{B}}(z)=\mathbb{E}[\Theta\,|\,Z=z]$. We then have the following, which is a re-statement of García-Trillos and Sen (2024, Theorem 2.4) and bears some similarity to Freirich et al. (2021, Theorem 1). Recall from Section 2 that $\Gamma(F;G)$ represents the space of all couplings of $F$ and $G$, and let us write $R_{\mathcal{DCB}}$ for the minimum value of problem ($\mathcal{DCB}$).
@@
$\underset{\pi\in\Gamma(F;G)}{\textnormal{minimize}}\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}(z,\eta)\,\textnormal{d}\pi(z,\eta),$ (3.4)
has a unique solution $\pi_{\mathcal{DCB}}$. This solution is concentrated on the graph of a function which we denote by $\delta_{\mathcal{DCB}}:\mathbb{R}^{d}\to\mathbb{R}^{m}$. This function $\delta_{\mathcal{DCB}}$ can be written as $\delta_{\mathcal{DCB}}=\nabla\phi\circ\delta_{\mathcal{B}}$ for some convex function $\phi:\mathbb{R}^{m}\to\mathbb{R}\cup\{+\infty\}$, and we have $R_{\mathcal{DCB}}=R_{\mathcal{B}}+W_{2}^{2}(G,(\delta_{\mathcal{B}})_{\#}F)$. Consequently, (3.4) is a tight convex relaxation of ($\mathcal{DCB}$), and $\delta_{\mathcal{DCB}}$ is the $F$-a.s. unique solution to ($\mathcal{DCB}$).
The above result says that, under appropriate conditions, problem ($\mathcal{DCB}$) indeed has a unique solution which can be obtained by solving the linear program (3.4). Further, the risk of the distribution-constrained denoiser $\delta_{\mathcal{DCB}}(Z)$ is the sum of the Bayes risk $R_{\mathcal{B}}$ and the squared Wasserstein distance between the distributions of $\Theta$ and $\delta_{\mathcal{B}}(Z)$. Contrast this with Theorem 3.3, where we showed that the risk of the variance-constrained denoiser $\delta_{\mathcal{VCB}}(Z)$ is sum of the Bayes risk and the squared Bures-Wasserstein distance between the covariance matrices of $\Theta$ and $\delta_{\mathcal{B}}(Z)$.
@@ 3.3 General-Constrained Denoising
Lastly, we consider problem ($\mathcal{GCB}$) of general-constrained Bayes denoising, which allows enough flexibility to encapsulate problems ($\mathcal{VCB}$) and ($\mathcal{DCB}$) as well as many others. Appendix C.1 contains our detailed discussion, including conditions under which problem ($\mathcal{GCB}$) admits a tight convex relaxation and a unique solution $\delta_{\mathcal{GCB}}$ (Theorem C.3), which is similar to the analogous result for problem ($\mathcal{DCB}$) (Theorem 3.8). As an example of a general-constrained denoising problem relevant to our application in Subsection 6.2 (see Appendix C for more examples), we consider the following:
@@
Suppose that $G$ is known to have support contained in a closed set $S\subseteq\mathbb{R}^{m}$. In that case, the posterior mean satisfies $\delta_{\mathcal{B}}(Z)\in S$ almost surely. However, the variance constrained denoiser $\delta_{\mathcal{VCB}}(\cdot)$ may take on values in $S$ with positive probability, which can be undesirable. Instead, one may aim to minimize the denoising risk subject to simultaneous constraints on the variance and the support of $\delta(Z)$, thereby solving
$\underset{\delta:\mathbb{R}^{d}\to\mathbb{R}^{m}}{\textnormal{minimize}}\,\,\mathbb{E}\left[\|\delta(Z)-\Theta\|^{2}\right]\quad\textnormal{s.t.}\quad\mathbb{E}[\delta(Z)]=\mathbb{E}[\Theta],\,\textnormal{Cov}(\delta(Z))=\textnormal{Cov}(\Theta),\,\,\mathbb{P}(\delta(Z)\in S)=1.$ (3.5)
To see that this can be cast as a special case of problem ($\mathcal{GCB}$), take $\psi(\eta):=\min_{\theta\in S}\|\eta-\theta\|$ and note that $\mathbb{P}(\delta(Z)\in S)=1$ is equivalent to $\mathbb{E}[\psi(\delta(Z))]=\mathbb{E}[\psi(\Theta)]=0$ In the case $m=d=1$ an important constraint on the support is nonnegativity, i.e., $S=[0,\infty)$.